OpenStudy (anonymous):

What is the solution to the following system of equations? 8x - 10y = -24 -8x + 6y = 40

OpenStudy (anonymous):

@campbell_st

OpenStudy (anonymous):

add the two equations simultaneously and you'll get -4y = 16. solve for y

OpenStudy (anonymous):

^8x cancels out there

OpenStudy (anonymous):

And the answer is...........

OpenStudy (anonymous):

Solve by elimination. The x terms can cancel immediately this way. 8x - 10y = -24 -8x + 6y = 40 -4y = 16 y = -16/4 If you want, you can also solve using substitution, but this is easier.

OpenStudy (anonymous):

Btw, you have to divide -16/4

OpenStudy (anonymous):

THen you plug it back in to solve for x and there is your solution set.

OpenStudy (anonymous):

You mean plug it back into y?

OpenStudy (anonymous):

Yeah. Plug what you got for y and put it in for y in one of the equations. 8x - 10(Put in value here) = -24 and solve for x.

OpenStudy (anonymous):

so x=-8 and y=-4?

OpenStudy (anonymous):

Yup :)

OpenStudy (anonymous):

Another way to write the solution is (-8, -4)